[Hallc_running] [New Logentry] RC Daily Update

cyero at jlab.org cyero at jlab.org
Sun Sep 18 08:40:01 EDT 2022

Logentry Text:
Last night, the Aluminum hole in target ladder was found by moving the target ladder 
by 4 mm downwards and setting the horizontal beam positions to the previous values used by pionLT run group. 
--> Hole Center  https://logbooks.jlab.org/entry/4043268
--> Beam Positions https://logbooks.jlab.org/entry/4043267

D. Meekins then came in and added corrected for ladder offset to all the targets, 
so now its OK to move to each target as intended (no need to calculations, etc.) 

Accelerator then proceeded to doing ion chamber calibrations, but had Ion Chamber functionality test failed
which prevented them from doing the actual ion chamber calibrations. The calibrations were suspended due to this test failure.
Systems Safety Group (SSG) is  investigating
---> https://logbooks.jlab.org/entry/4043301
---> https://logbooks.jlab.org/entry/4043331

After speaking with D. Gaskell this morning, it is OK with him for Ops to bypass the Compton Polarimeter Ion Chamber
if needed, as a temporary solution. This possibility will be discussed at the MCC  9AM meeting today.

Another possibility is for the Ion Chamber calibrations t to be done with tune beam rather than CW, but
that will require some additional re-calibration by MCC. Will also be discussed at meeting MCC meeting.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4043409
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220918/88839373/attachment.html>

More information about the Hallc_running mailing list